1. Find scalars a,b,c that are real numbers such that at least one of a,b,c is non zero:

at] +b]+cE]=R]

2. Find a nonzero vector v in R^4 orthogonal to:

-3
Answer #1

Tatb+3c7 a -b - C a tb+ 36=o a-b = C. a-b -C =0 from O þut c= a -b a +b+ 3 Ca-b) = 0 arb +39 -3b=0 44 -2b =0 : C=a -b El-E 2a-3 92 -2 3 + =D0 in particulur teuke x> =Q3=dy =1 → di= 6+4 -2 %3D Hance nozero vechar 28/27 X2 o her Cases alio exist